DanielRichter:

Complete the square to rewrite the following equation. Identify the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 20

AZ:

When you're completing the square, you're trying to make it in the form of a perfect square which is \( a^2 + 2ab + b^2 = (a+b)^2\) or \( a^2 -2ab + b^2 = (a-b)^2 \)

DanielRichter:

Ok

AZ:

So for your question, we have \( (x^2 + 2x +\color{red}{ \text{__}}) + (y^2 + 2y +\color{orange}{ \text{__}}) = 20 + \color{red}{\text{__}}+ \color{orange}{ \text{__}}\) so we're trying to find a number to add to \(x^2 + 2x\) so that way we can simplify it into something like (a+b)^2 and the same thing goes for \(y^2 + 4y\) and to make sure that we're not changing the equation, we have to add both of those values to the other side of the equal sign as well Does that make sense so far?

DanielRichter:

Yes

AZ:

Excellent, so first let's find what we have to add to x^2 + 4x \( a^2 + 2ab + b^2 = (a+b)^2\) \( x^2 + 2x + \color{red}{\text{__}} = (x + ?)^2\) Do you see how 'a' and 'x' are the same thing? So we're trying to find that third term, but to do that we need to know what 'b' is The middle term in the formula is 2ab And in your question, we have 2x in the middle We said that our 'a' is 'x'. Can you figure out what 'b' would have to be?

DanielRichter:

y

DanielRichter:

'B' = 'Y '

AZ:

No, we're trying to get your question to look like the formula so that way we can simplify it We want the terms to be equal 2ab = 2x and since we said that 'a' and 'x' are the same (We can easily tell this because both start with a^2 and x^2) so we get 2xb = 2x what is the value of 'b'

DanielRichter:

is it 0

AZ:

No What does 'b' have to be for 2xb to be equal to 2x Anything multiplied with 0 is going to be 0 2x * 0 = 0 2x * b = 2x what is b? if you divide 2x on both sides, what do you get?

DanielRichter:

1

DanielRichter:

b=1

AZ:

Exactly!! b = 1 and so that last term is b^2 if you look in the formula, we're adding by b^2 so that way we can simplify it to (a+b)^2 so what is 1^2 = ?

DanielRichter:

1

AZ:

Exactly! So now we have \( a^2 + 2ab + b^2 = (a+b)^2\) \( x^2 + 2x + \color{red}{\text{1}} = (x + 1)^2\) Do you see how we did that? We found out that b was 1, so we had to add b^2 and then we could simplify it into \((a+b)^2\)

AZ:

So now we have \( (x^2 + 2x +\color{red}{ \text{1}}) + (y^2 + 2y +\color{orange}{ \text{__}}) = 20 + \color{red}{\text{1}}+ \color{orange}{ \text{__}}\) \( (x+1)^2~~~~~~~~~ + (y^2 + 2y +\color{orange}{ \text{__}}) = 20 + \color{red}{\text{1}}+ \color{orange}{ \text{__}}\) Do you follow?

DanielRichter:

yes

DanielRichter:

i Do

AZ:

So we just have to figure out what we add to \(y^2 + 4y\) and then we'll be all done I apologize for my typo in my previous comments where I wrote `2y` Your question has `4y` So it's the same concept \( a^2 + 2ab + b^2 = (a+b)^2\) \( y^2 + 4y + \color{orange}{\text{__}} = (y + ?)^2\) The middle term in the formula is 2ab And in your question, we have 4y in the middle Since the 'a' is 'y' can you figure out what 'b' would have to be? 2ab = 4y a is the same y 2yb = 4b what does b have to be?

DanielRichter:

2

AZ:

Perfect, so what is b^2 and what can we simplify it all down to?

DanielRichter:

4

DanielRichter:

is it 4

AZ:

yes, b^2 is 4 so we're adding 4 and remember that it gets simplified down to (a+b)^2 so 'b' is the term when we factor it

DanielRichter:

is it x2

AZ:

It's just 2

AZ:

Because \( a^2 + 2ab + \color{orange}{b}^2 = (a + \color{orange}{b})^2 \) and we found that b = 2 so \( y^2 + 4y + \color{orange}{2}^2 = (a + \color{orange}{2})^2\) Does that make sense?

DanielRichter:

yes

AZ:

But remember, we're adding 2^2 on both sides of the equation!!! And the equation gets simplified to (y+2)^2 \( (x+1)^2~~~~~~~~~ + (y^2 + 2y +\color{orange}{ \text{4}}) = 20 + \color{red}{\text{1}}+ \color{orange}{ \text{4}}\) Do you see what I mean? Since we added 4 to that equation, we have to add 4 to the other side and then to complete it, we just have to factor down the \(y^2 + 4y + 4\) into \( (y+2)^2\) and just add the numbers on the other side of the equal sign Can you do that?

DanielRichter:

y^2+ 4y +8

AZ:

no?? you just need to replace \(y^2 + 4y + 4\) with \( (y+2)^2\) remember all that work we did earlier??

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Because \( a^2 + 2ab + \color{orange}{b}^2 = (a + \color{orange}{b})^2 \) and we found that b = 2 so \( y^2 + 4y + \color{orange}{2}^2 = (y+ \color{orange}{2})^2\) Does that make sense? \(\color{#0cbb34}{\text{End of Quote}}\)

AZ:

and then you just have to add up 20 + 1 + 4 = ??

DanielRichter:

Oh ok so it is 5

AZ:

are you sure?

DanielRichter:

Thanks a lot for the help I really appreciate it.

DanielRichter:

oh I mean 25

AZ:

You're most welcome!! We haven't identified the center and the radius just yet, however

DanielRichter:

sorry about the 5 thing

DanielRichter:

ok

AZ:

So now we have the equation in the form of \( (x+1)^2+ (y+2)^2 = 25\) And this is the standard form of a circle \( (x-h)^2 + (y-k)^2 = r^2\) (h, k) is going to be the center of your circle and r is going to be the radius

DanielRichter:

ok

DanielRichter:

So how does one solve it

AZ:

The numbers are all right there in the same form

AZ:

Here's some colors to help you see it \( (x+1)^2+ (y+2)^2 = \color{red}{25}\) And this is the standard form of a circle \( (x-h)^2 + (y-k)^2 = \color{red}{r^2}\) (h, k) is going to be the center of your circle r is your radius so what is the radius? if r^2 is equal to 25 what is r equal to??

DanielRichter:

5

AZ:

Good, so that's your radius Now what about the center of your circle \( (x+1)^2+ (y+2)^2 = 25\) \( (x-h)^2 + (y-k)^2 = r^2\) (h, k) is going to be the center of your circle be careful about the signs we have (x-h) there's a minus but the x-value of the center of the circle is 'h' so in (x+1) you have to flip the sign, the x-value isn't going to be 1 but -1 since it's x PLUS 1 and not x - 1 does that make sense?

DanielRichter:

somewhat

AZ:

so what is the center? I gave you the 'h' which is the x-value of the center what is the 'k' do the same thing that's your y-value of the center

AZ:

An image that might help you understand what it means by that (h, k) is the center

1 attachment
DanielRichter:

It is 4

DanielRichter:

but probably not

AZ:

How? Look at the numbers

AZ:

\( (x\color{red}{+1})^2+ (y\color{orange}{+2})^2 = 25\) \( (x \color{red}{-h})^2 + (y\color{orange}{-k})^2 = r^2\)

AZ:

so what is 'h' and what is 'k'

AZ:

what does 'h' have to be to make x-h into x+1 what does 'k' have to be to make y-k into y+2

DanielRichter:

2 3

AZ:

uhhhhh those aren't even the numbers that's listed?? All you have to do is flip the sign

DanielRichter:

what I first thought but it did not make sense mutch though so, -1 -2

AZ:

so that's the h and k and you write the center as (h, k)

DanielRichter:

(-1,-2)

DanielRichter:

is it that

AZ:

Exactly!

DanielRichter:

so is it 5 and (-1,-2)

AZ:

yes